# PerspectiveNet: Multi-View Perception for Dynamic Scene Understanding

### v1

@@ Abstract
Generating detailed descriptions for events from multiple cameras and viewpoints is challenging due to the complex and inconsistent nature of the visual data. In this paper, we present PerspectiveNet, a lightweight yet efficient model to generate long descriptions for multiple-view cameras by using a vision encoder, a small connector module to convert visual features to a fixed-size tensor, and large language models (LLM) to utilize the strong capability of LLM in natural language generation. The connector module is constructed with three primary objectives, map visual features onto LLM embedding, emphasize the critical information needed for generating descriptions, and yield the fixed-size feature matrix as the output. Furthermore, we augment the prowess of our solution by employing a second task, the correct sequence of frame detection, to enable the model’s ability to search for the correct frame sequence to generate the description. Finally, we combine the connector module, a second training task, a large language model, and a visual feature extraction model into a single model and train it for the Traffic Safety Description and Analysis task, in which the model has to generate long fine-grain descriptions for an event from multiple cameras with and viewpoints. The resulting model is lightweight and ensures an efficient training and testing process, yet very effective.
@@ 1 Introduction
In recent years, there has been a variety of research about large language models (LLMs), those research has improved the capability of LLMs in natural language understanding, reasoning, knowledge updating, and instructions following. Moreover, LLMs can be adapted to different domains by fine-tuning different datasets, which shows their potential to assist people, enhance productivity, and handle many difficult problems. However, since LLMs are just about text and can only work with language tasks, they cannot understand different information sources such as sound and vision, which are crucial to solving many real-world problems. For those reasons, there is much research about integrating LLMs with vision models to enable LLMs with visual tasks such as LLaVA, Qwen-VL, InstructBLIP, and Flamingo. These works have advanced the capability of LLMs in visual problems significantly.
While problems with images and videos are the main target of many vision language models, multiple viewpoints of an event are not on the priority list. As a result, most models are not effectively designed and trained to handle such visual information. Different from video data and images, which only show objects and events from one viewpoint, this type of data can provide more information about an event by inspecting from different positions and help vision language models understand and generate descriptions about the event more precisely. While the number of videos and frame number of each video can be various, handling such information can be more challenging than videos and images. This is because the large and undefined number of videos results in the large and undefined number of features, which leads to the difficulty in finding the similarity of relation between many events, and frames in videos. Moreover, since the number of features is large, putting all of them into the LLMs to handle might not be the best choice due to the excessive increase in size of the input, and the increase of computation cost. To solve this, visual features should be encoded to smaller sizes but still need to be informative for language models to generate long fine-grain descriptions.
In this paper, with the awareness of those problems, we introduce PerspectiveNet, a method to connect a large and undefined number of visual features with LLMs which can be applied to many visual language tasks. The module is constructed to take visual features of multiple videos and viewpoints as input and create the output with 3 features: small, informative, and size-consistent to provide context for LLMs to generate. Moreover, we also present a second task during the training process to increase the capability of our module in detecting important events and frames. Finally, we experiment by integrating the above methods on multiple videos and viewpoints datasets with the main task being traffic safety description and analysis.
@@
  • Create a module to connect the visual features of multiple cameras and viewpoints with LLMs.
  • Propose a second task during the training process to help the model detect important features from many videos and ignore unimportant ones.
  • We experiment with those methods in a multiple viewpoints dataset and report our result.
@@ 2 Related work
Recently, with the release of many large language models like Mistral[10], LLaMA[17], Bloom[11], Phi[6], or Qwen[2], become popular due to the superior performance in natural language understanding and generation. Moreover, they have the potential to adapt to different domains and tasks, not only in natural language but also in different domains like computer vision. Many research groups have proposed solutions to integrate vision models with large language models to create vision language models (VLMs) for many tasks like visual question-answering or caption generation. One of the common solutions to create VLMs is to encode vision information (images, videos) to a feature tensor, slightly modify that feature by changing shape, and information with a module, and then feed that matrix to LLMs through cross-attention layers or by using these tensor as input context for LLM to generate the answer.
Many previous works to create VLMs follow this general structure such as LLaVA[15], Qwen-VL[2], and InstructBLIP[4] that used vision encoder to encode the image and get visual features, then feed those features to a small module before using the output as input context for large language model to generate captions, descriptions and answers. Meanwhile, some vision language models like Flamingo[1], mPLUG[12] or BLIP[13] decided to use some more complex solutions to modify the visual features before feeding to decoder-only through cross attention layers to generate natural text. Most vision language models use Vision Transformer (ViT)[5] as a visual encoder due to the strong capability of this architecture in extracting features from visual information despite its computation cost.
Moreover, previous research has presented different methods to connect visual information to language models. A noticeable contribution is using cross attention[1] to convert visual features of different sizes to a constant size tensor before being fed to a language model to generate text. Another important research is Q-former[14], this method and module can teach itself to focus on visual information of an image which is the most relevant to the text. Most of those works focus on improving the ability of the connector. By carefully selecting and modifying visual features, the connector can create a new visual matrix that is informative and precise and reduces the noise of non-important information in images and videos. Furthermore, if the connector can return a small output while keeping the most important part of visual information for the language model to understand and generate, this could significantly reduce the computational cost during the training or inference progress. Not only changing the structure of the connector but recent research also pays attention to training models on multiple objectives[13] with some tasks like ITM (Image-Text Matching loss), ITC (Image-Text Contrastive loss), LM (Language model loss). This strategy shows great improvement for VLMs since it can improve the comprehension of the model on the relation of those objectives. Lastly, instruction learning has improved the ability of many LLMs (InstructGPT[16], PLAN-PaLM[3], and OPT-IML[8]) also show a process on VLM[15].
However, while there is much research about using VLMs on visual data like images, and video, types of data like multiview-camera do not attract significant attention even though this type of data can provide more detail by inspecting an event from multiple positions. As a result, being inspired by many previous works on VLMs and their limitations on such data, PerceptiveNet is presented as a way to construct such VLM for this specific data type.
@@ 3 Approach
In this session, we start with a brief overview of our description generation task, and the dataset on which we train and test our model. After that, we introduce the details of our solution for this problem, which includes the overall architecture of the PerceptiveNet, more details of our vision language connector, and the training method.
@@ 3.1 Task and dataset
This paper focuses on generating long fine-grained video descriptions for many traffic safety scenarios and pedestrian accidents. The resulting model should be able to describe the continuous moment before the incidents and normal scenes, as well as all the details about the surrounding context, attention, location, pedestrians, and vehicles. The scene of events might be recorded from different viewpoints and positions, this requires the model to understand the similarity between each viewpoint and the relevance of many frames in an event to generate informative descriptions.
The dataset used in this paper is the WTS, a newly created dataset. Each sample of the dataset includes 2 long and detailed descriptions for vehicle and pedestrian, $n$ videos ($n\geq 1$) about a sequence of events about traffic, incident on the street, the start time and end time of the event that the model should focus on to describe, annotated segment (pre-recognition, recognition, judgment, action, and avoidance), and bounding boxes of the instance (pedestrian or vehicle) that relates to the description. There are two types of bounding boxes generated (machine) and annotated (human).
To handle this data better, we refactor the dataset structure. After this data pre-processing step, each sample of the dataset includes $n$ videos denoted as $V=[v_{0},v_{1},\dots,v_{n-1}]$, the frame at second $s$ of video $v_{i}$ is denoted as $v_{is}$, one long description $X=[x_{0},x_{1},...x_{t-1}]$ with $t$ is the length of description, a target object to describe $o$ (pedestrian or vehicle), start time $m$, end time $n$, segment $s$ (pre-recognition, recognition, judgment, action, and avoidance). In general, the task can be described as follows:
@@
Furthermore, it is worth mentioning that our solution does not use any information from given bounding boxes, which explains why the post-processing dataset does not contain any information about bounding boxes, and even when there are many videos about an event, all start simultaneously. For that reason, at the same time, the frames of all cameras show the same scene from different positions and the visual context of those frames should be relatively similar. Besides multiple-view cameras, there are data samples with only one video, which makes the task for those samples become video description generation.
@@ 3.2 Architechture
Figure 1: Architecture overview
@@
  • Visual encoder: to get visual information from each frame of each video, we use the pre-trained image encoder of BLIP-2[14] (ViT-g/14).
  • Vision-Language Connector: for the connector, we use two Perceiver module[9]. The first Perceiver is used to convert features of all frames of different videos at the same time to a single feature vector. The second Perceiver is employed to reduce the shape of the long video feature to a smaller constant shape.
  • Large language model: we avoid using LLMs with more than 3 billion parameters for the generating task due to the limited resources. Finally, we decided to use Phi-1.5[6] due to its small size and strong capability in natural language generation despite the model having not been trained on any visual information yet. Moreover, we also employ LoRA[7] due to the need to fine-tune LLMs on the new task.
For each data sample, we are provided with $n$ videos, where we extract frames from those videos at a frequency of 1 frame per second. It is important to note that even though all videos record the same event, not all of them have the same length, some might finish sooner than expected.
@@
Phi 1.5: The new visual information is then used as a context token for Phi 1.5 to generate a description. The initial input, which provides information for LLM, includes the visual tokens, target object (pedestrian or vehicle), and segment (pre-recognition, recognition, judgment, action, or avoidance). Moreover, to make the language model adapt well to this task, LoRA[7] layers are injected into the attention mechanism of Phi 1.5. While language model weight is frozen during training, weights of the LoRA adapter are still set to trainable. This step helps the language model to understand its task better after our training progress while maintaining a strong ability in text generation.
Cross-attention layers can be added to the language model to receive visual information, however, adding this module would increase the size of the language model significantly. Furthermore, since the newly added layers are initial with random weight, those layers will also need to be trained during the training process, which would highly increase the VRAM consumption. As a result, using visual information as extra visual context in the input prompt would be the best choice for limited resources.
@@ 3.3 Training Strategy
In each data sample, we are provided with the start time and end time of the event that needs to be described. Denote that input videos is $V=[v_{0},v_{1},\dots,v_{t-1}]$ where $v_{i}$ is all frames at time $i$, with that notation, the event from start time $m$ to end time $n$ can be written as $E=[v_{m},v_{m+1},\dots,v_{n}]$. With two visual inputs and the requirement to generate the same description, we fed both of them to the Visual Encoder and then the adapter to collect $f_{V}$ and $f_{E}$, which are the final visual context of full video and partial events respectively. Then, the dot product of $f_{V}$ and $f_{E}$ is calculated and compared with $I$ (Identity matrix) by cross-entropy loss.
@@
Our main target for this loss function is to provide the same context but a different part of videos to our connector and train the connector to understand where the important event should be focused on to generate the context. While the $f_{E}$ is the final feature of the correct event, the $f_{V}$ after training should ignore every frame but frames can provide information about the event that needs to be described.
Figure 2: Large language model input and output
For the generation task, the prediction of visual tokens is the token at index 0 (which is ‘!’). However, the last visual token should predict the first text token in the prompt as shown in Fig. 2. Besides adding extra visual tokens, this model performs like a normal large language model.
@@
Finally, the total loss is calculated by adding the matching loss $E_{M}$ and generation loss $E_{G}$. By taking the backpropagation with this loss, the model can perform updating on both generation and event matching task
@@ 4 Experiment
Since our vision encoder is frozen and not being trained during the training process, we decided to extract visual features of videos, save them to files, and use those features in our training process instead of re-running the whole visual encoder every time. By doing this we preserve more VRAM for training LLM and the adapter. For the experiment, both the training and testing processes are done with the hardware including 30GB RAM, and GPU P100 with 16GB VRAM, which can easily be found and used on Kaggle.
Our model was trained on the WTS dataset two times. For the first time, we train our model with the WTS dataset except for the external and normal trimmed data. The training configuration includes 20 epochs, the learning rate is 5e-5, we use Adam optimizer, the batch size is 2, the evaluate step is 100, and the gradient accumulation step is 2. The result of this training process is shown in 1. Even though the model was trained on a small internal dataset and is predicted to perform poorly on the external dataset, the result shows the opposite when the model can achieve a high score, nearly 24, on the new dataset with a nearly different context. However, this result on the external dataset is still far behind the experiment result on the internal dataset.

### v2

# PerspectiveNet: Understand Multi-View Videos with Event Matching for Traffic Safety Description and Analysis

@@ Abstract
Generating detailed descriptions for events from multiple cameras and viewpoints is challenging due to the complex and inconsistent nature of the visual data. In this paper, we present PerspectiveNet, a lightweight yet efficient model to generate long descriptions for multiple-view cameras by using a vision encoder, a small connector module to convert visual features to a fixed-size tensor, and large language models (LLM) to utilize the strong capability of LLM in natural language generation. The connector module has three primary objectives: map visual features onto LLM embedding, emphasize the critical information needed for generating descriptions, and yield the fixed-size feature matrix as the output. Furthermore, we enhance the model’s ability by incorporating a second task, sequence frame detection, which enables the model to search for the relevant frame sequence. Finally, we combine the connector module, a second training task, a large language model, and a visual feature extraction model into a single model and train it for the Traffic Safety Description and Analysis task, in which the model has to generate long fine-grain descriptions for an event from multiple cameras with and viewpoints. The resulting model is lightweight and efficient for both training and inference.
PerspectiveNet: Understand Multi-View Videos with Event Matching for Traffic Safety Description and Analysis
Phu-Vinh Nguyen Uppsala University phu-vinh.nguyen.1216@student.uu.se                        Tan-Hanh Pham Florida Institute of Technology, USA hanhpt.phamtan@gmail.com
@@ 1 Introduction
In recent years, there has been a variety of research about large language models, improving the capabilities of LLMs in natural language understanding, reasoning, knowledge updating, and instruction following. Moreover, LLMs can be adapted to different domains by fine-tuning different datasets, which shows their potential to assist people, enhance productivity, and handle many difficult problems. However, since LLMs are just about text and can only work with language tasks, they cannot understand different information sources such as sound and vision, which are crucial to solving many real-world problems. For those reasons, there is much research about integrating LLMs with vision models to enable LLMs with visual tasks such as LLaVA Liu et al. (2023), Qwen-VL Bai et al. (2023b), InstructBLIP Dai et al. (2024), and Flamingo Alayrac et al. (2022), which advances the LLMs’ capability in visual tasks significantly.
While problems with images and videos are the main target of many vision language models, multiple viewpoints of an event are not on the priority list. As a result, most models are not effectively designed and trained to handle such visual information. Different from video data and images, which only show objects and events from one viewpoint, this type of data can provide more information about an event by inspecting it from different positions and help vision language models understand and generate descriptions about the event more precisely. While the number of videos and the frame number of each video can be various, handling such information can be more challenging than videos and images. This is because the large and undefined number of videos results in a large and undefined number of features, which leads to the difficulty in finding the similarity of relation between many events and frames in videos. Moreover, since the number of features is large, putting all of them into the LLMs to handle might not be the best choice due to the excessive increase in size of the input and the increase in computation cost. To solve this, visual features should be encoded to smaller sizes, but still need to be informative for language models to generate long, fine-grain descriptions.
In this paper, with the awareness of those problems, we introduce PerspectiveNet, a method to connect a large and undefined number of visual features with LLMs, which can be applied to many visual language tasks. The module is constructed to take visual features of multiple videos and viewpoints as input and create the output with 3 features: small, informative, and size-consistent to provide context for LLMs to generate. Moreover, we also present a second task during the training process to increase the capability of our module in detecting important events and frames. Finally, we experiment by integrating the above methods on multiple videos and viewpoints datasets, with the main task being traffic safety description and analysis.
@@
  • Create a module to connect the visual features of multiple cameras with LLMs.
  • Propose a second task during the training process to help the model detect important features and ignore unimportant ones.
  • We experiment with those methods in a multiple viewpoints dataset and report our results.
@@ 2 Related work
In recent years, large language models (LLMs) such as Mistral Jiang et al. (2023), LLaMA Touvron et al. (2023), Bloom Le Scao et al. (2022), Phi Gunasekar et al. (2023), and Qwen Bai et al. (2023a) have shown exceptional performance across various natural language processing tasks. Their success has encouraged the development of vision-language models (VLMs) that combine visual inputs with LLMs to solve tasks like visual question answering and caption generation. These VLMs generally use a vision encoder to convert visual data into feature embeddings, which are then passed to the LLM via cross-attention or adapted as contextual tokens to generate responses.
Many well-known VLMs such as LLaVA Liu et al. (2024), Qwen-VL Bai et al. (2023a), and InstructBLIP Dai et al. (2024) follow a modular approach by combining pretrained vision encoders (e.g., ViT Dosovitskiy et al. (2020)) with LLMs through lightweight projection layers or cross-attention modules. More complex architectures like Flamingo Alayrac et al. (2022), mPLUG Li et al. (2022a), and BLIP Li et al. (2022b) integrate visual features using sophisticated mechanisms such as gated cross-modal transformers or multi-level attention layers. While these methods improve reasoning and text generation performance, they often require large memory footprints (e.g., LLaVA-34B, Qwen-VL-Chat-14B) and multi-GPU infrastructure, which limits their deployment in real-world applications or resource-constrained environments.
To bridge the gap between performance and efficiency, research has explored the design of more effective "connectors" that transform visual inputs into compact yet informative representations for LLMs. One popular method is Q-Former Li et al. (2023), which learns to query and distill relevant visual information from high-dimensional image features. Other works like Flamingo Alayrac et al. (2022) employ cross-attention layers to align multi-modal sequences of different lengths. Additionally, multi-task learning objectives such as image-text matching (ITM), image-text contrastive learning (ITC), and language modeling (LM) have been incorporated to guide training and improve alignment across modalities Li et al. (2022b). Recent studies also show that instruction tuning (e.g., InstructGPT Ouyang et al. (2022), OPT-IML Iyer et al. (2022)) significantly enhances the zero-shot and few-shot performance of VLMs.
Although some existing methods support VLMs to understand multi-camera information and have promising results, like CityLLaVA Duan et al. (2024), Divide&Conquer Xuan et al. (2024), those models still require a strong VLM baseline to produce good results. In contrast, our work introduces PerspectiveNet, a lightweight vision-language architecture tailored for multi-camera pedestrian-centered video captioning. Without relying on prior pretraining from large-scale VLMs, our method achieves competitive performance on the WTS dataset Kong et al. (2024) while maintaining a smaller size and reduced computational cost. Our method shows significantly better performance compared to methods with a similar size, like Multi-perspective+Refinement To et al. (2024).
@@ 3 Approach
In this session, we start with a brief overview of our description generation task and the dataset on which we train and test our model. After that, we introduce the details of our solution for this problem, which includes the overall architecture of the PerceptiveNet, more details of our vision language connector, and the training method.
@@ 3.1 Task and dataset
This paper focuses on generating long, fine-grained video descriptions for many traffic safety scenarios and pedestrian accidents. The resulting model should be able to describe the continuous moment before the incidents and normal scenes, as well as all the details about the surrounding context, attention, location, pedestrians, and vehicles. The scene of events might be recorded from different viewpoints and positions, which requires the model to understand the similarity between each viewpoint and the relevance of many frames in an event to generate informative descriptions.
The dataset used in this paper is the WTS Kong et al. (2024). Each sample of the dataset includes 2 long and detailed descriptions for vehicle and pedestrian, $n$ videos from different viewpoints ($n\geq 1$) about a sequence of events about traffic, incident on the street, the start time and end time of the event that the model should focus on to describe, annotated segment (pre-recognition, recognition, judgment, action, and avoidance), and bounding boxes of the instance (pedestrian or vehicle) that relates to the description.
To handle this data better, we refactored the dataset structure. After this data pre-processing step, each sample of the dataset includes $n$ videos denoted as $V=[v_{0},v_{1},\dots,v_{n-1}]$, the frame at second $s$ of video $v_{i}$ is denoted as $v_{is}$, one long description $X=[x_{0},x_{1},...x_{t-1}]$ with $t$ is the length of description, a target object to describe $o$ (pedestrian or vehicle), start time $m$, end time $n$, segment $s$ (pre-recognition, recognition, judgment, action, and avoidance). In general, the task can be described as follows:
@@
Furthermore, it is worth mentioning that our solution does not use any information from the given bounding boxes, which explains why the post-processing dataset does not contain any information about bounding boxes, and even when there are many videos about an event, they all start simultaneously. For that reason, at the same time, the frames of all cameras show the same scene from different positions, and the visual context of those frames should be relatively similar. Besides multiple-view cameras, there are data samples with only one video, which makes the task for those special samples video description generation.
@@ 3.2 Architechture
Figure 1: Architecture overview, videos are lists of frames or images. Each frame will be fed to the vision encoder to extract a list of embeddings for each video. After that, the Perceiver adapter is applied to all frames at the same moment to get the multi-view embedding. Next, it is concatenated with the task embedding before feeding it to another Perceiver layer and LLM to get the description of the event. In this image, gray color denotes input and output data, and blue denotes the frozen modules, which are not trained. Red denotes trained modules during the training phases. Finally, yellow denotes the output and input of modules.
@@
  • Visual encoder: to get visual information from each frame of each video, we use the pre-trained image encoder of BLIP-2Li et al. (2023) (ViT-g/14).
  • Vision-Language Connector: for the connector, we use two Perceiver moduleJaegle et al. (2021). The first Perceiver is used to convert features of all frames of different videos at the same time to a single feature vector. The second Perceiver is employed to reduce the shape of the long video feature to a smaller constant shape. By doing this, we only extract important visual information and use LLM to generate descriptions. This reduces the redundancy and computation cost, ensuring the input visual context of LLM is not too long.
  • Large language model: We avoid using LLMs with more than 3 billion parameters for the generating task due to the limited resources. Finally, we decided to use Phi-1.5Gunasekar et al. (2023) due to its small size and strong capability in natural language generation, despite the model not having been trained on any visual information yet. Moreover, we employ LoRAHu et al. (2021) due to the need to fine-tune LLMs on the new task.
For each data sample, we are provided with $n$ videos, from which we extract frames from those videos at a frequency of 1 frame per second. It is important to note that even though all videos record the same event and start at the same time, not all of them have the same length; some might finish sooner than expected.
@@
Task Embedding: To generalize the model’s ability to generate descriptions at each specific part of the event, we use an embedding layer to embed each task. This information is important for the model to understand what information, object, or part of the event should be focused on. In the experiment section, our dataset has five tasks and two main objects (human and vehicle), making the embedding contain ten tokens to represent all cases.
Phi 1.5: The new visual information is then used as a context token for Phi 1.5 to generate a description. The initial input, which provides information for LLM, includes the visual tokens, target object (pedestrian or vehicle), and segment (pre-recognition, recognition, judgment, action, or avoidance). Moreover, to make the language model adapt well to this task, LoRAHu et al. (2021) layers are injected into the attention mechanism of Phi 1.5. While the language model weight is frozen during training, the weights of the LoRA adapter are still set to trainable as shown in fig. 2. This step helps the language model to understand its task better after our training progress, while maintaining a strong ability in text generation.
Cross-attention layers can be added to the language model to receive visual information; however, adding this module would increase the size of the language model significantly. Furthermore, since the newly added layers are initialized with random weights, those layers will also need to be trained during the training process, which would highly increase the VRAM consumption. As a result, using visual information as extra visual context in the input prompt would be the best choice for our limited resources.
@@ 3.3 Training Strategy
In each data sample, we are provided with the event’s start and end times that need to be described. Denote that input videos is $V=[v_{0},v_{1},\dots,v_{t-1}]$ where $v_{i}$ is all frames at time $i$, with that notation, the event from start time $m$ to end time $n$ can be written as $E=[v_{m},v_{m+1},\dots,v_{n}]$. With two visual inputs and the requirement to generate the same description, we fed both of them to the Visual Encoder and then the adapter to collect $f_{V}$ and $f_{E}$, which are the final visual context of full video and partial events, respectively. Then, the dot product of $f_{V}$ and $f_{E}$ is calculated and compared with $I$ (Identity matrix) by cross-entropy loss.
@@
Our main target for this loss function is to provide the same context (event and object to describe) but a different part of the videos (all videos and just the target event in the video) to our connector and train the connector to understand where the important event should be to generate the context. While the $f_{E}$ is the final feature of the correct event, the $f_{V}$ after training should ignore every frame, but frames can provide information about the event that needs to be described. Furthermore, the idea of using the Cross-Entropy loss in this scenario is that we want each embedding vector in the matrix should be distinguished. This means that given two random tokens in the visual embedding at different positions, their information and embeddings must be different. This ensures that the visual information will be diverse and spread across all scenes without focusing on a small part of the videos and becoming overfitted.
Figure 2: Training on the large language model, the model’s input is only the visual tokens from previous steps. LoRa is employed to reduce the number of training parameters and to improve the model’s performance on the new task. In this image, gray color denotes input and output data, and blue denotes the frozen modules, which are not trained. Red denotes trained modules during the training phases. Yellow denotes the output and input of modules. Finally, orange is the lowest value that we need to reduce.
Figure 3: Second loss function on visual information to train the model to focus on important frames. The two inputs (visual tokens) are embeddings of the entire multi-view videos and of the targeted event. In this image, gray color denotes input and output data, and blue denotes the frozen modules, which are not trained. Yellow denotes the output and input of modules. Finally, orange is the loss value that we need to reduce.
For the generation task, the prediction of visual tokens is the token with index 0. To train this model, we use LoRA to reduce the number of training parameters as shown in fig. 2. Besides using the special token for the output of visual inputs, the remaining training process, including computing loss, remained the same.
@@
Finally, the total loss is calculated by adding the matching loss $E_{M}$ and the generation loss $E_{G}$. By taking the backpropagation with this loss, the model can perform updates on both the generation and event matching tasks
@@ 4 Experiment
4.1 Dataset description
Figure 4: Example of the multi-view video in the dataset WTS at a specific time. All video in the dataset starts at the same time. However, the length of each video in a set might vary. In this sample, the correct pedestrian behavior caption is: The pedestrian found himself positioned directly in front of the assailant vehicle with their orientation opposite to it. There was no relative distance; they were almost touching the vehicle. His line of sight remained focused on the vehicle, closely watching its actions. The pedestrian was about to stand still after moving at a slightly higher speed, and they were decelerating. The image is collected from the original paper of WTS
The Woven Traffic Safety (WTS) dataset, introduced in Track 2 of the AICity Challenge 2024, is designed for fine-grained understanding of pedestrian-vehicle interactions in traffic scenarios. It contains 155 multi-view scenarios (810 videos in total), recorded at 1080p and 30 fps from overhead and vehicle-mounted cameras. Each scenario includes five temporal phases: pre-recognition, recognition, judgment, action, and avoidance. Each segment of a scenario includes two detailed natural language captions, one for pedestrian behavior and one for vehicle behavior, averaging around 59 words and based on a structured checklist of over 170 items. Furthermore, to support structured reasoning, the dataset includes approximately 180 VQA items and over 3,000 additional annotated videos from BDD100K. Caption evaluation uses LLMScorer, a semantic similarity metric based on large language models, which captures fine-grained alignment between generated and reference descriptions. An example sample can be seen at fig. 4.
4.2 Evaluation Results
Since our vision encoder is frozen and not being trained during the training process, we decided to extract visual features of videos, save them to files, and use those features in our training process instead of re-running the whole visual encoder every time. By doing this, we preserve more VRAM for training the LLM and the adapter. For the experiment, both the training and testing processes are done with the hardware, including 30GB RAM and a GPU P100 with 16GB VRAM.
Our model was trained on the WTS dataset twice. In the first phase, we trained the model on an internal dataset, which represents a small subset of the full dataset. The training configuration included 20 epochs, a learning rate of 5e-5, the Adam optimizer, a batch size of 2, an evaluation step of 100, and a gradient accumulation step of 2. The results of this training phase are presented in table 1. Although the model was trained on a limited internal dataset and was expected to perform poorly on the external dataset, the results suggest otherwise. The model achieved a high score, nearly 24, on a new dataset with a different context. However, its performance on the external dataset still lags behind the results on the internal dataset. The strong performance after the initial training indicates the robustness of the architecture. Despite being trained on a smaller dataset and for more epochs, the model did not overfit and maintained strong generalization on the external test set. This behavior may be attributed to the event matching component and the diverse information captured during training $\mathcal{L}_{M}$.